# On the presence of a fifth force at the Galactic Center

### v3

@@ Abstract
Aims. The presence of a Yukawa-like correction to Newtonian gravity is investigated at the Galactic Center, leading to a new upper limit for the intensity of such a correction.
Methods. We perform a Markov Chain Monte Carlo analysis using the astrometric and spectroscopic data of star S2222 collected at the Very Large Telescope by GRAVITY, NACO and SINFONI instruments, covering the period from $1992$ to $2022$.
Results. The precision of the GRAVITY instrument allows us to derive the most stringent upper limit at the Galactic Center for the intensity of the Yukawa contribution ($\propto\,\alpha e^{-\lambda r}$) to be $|\alpha|<0.003$ for a scale length $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,\rm AU)$. This improves by roughly one order of magnitude all estimates obtained in previous works.
@@ 1 Introduction
General Relativity (GR) is the current, successful theory of gravity whose predictions have been extensively tested at Solar System scales and with gravitational waves emission by black holes (BHs) and binary pulsars (Will, 2014, 2018b; Nitz et al., 2021). Until now, no significant deviation from GR has been detected in any of these observations. However, it is also known that beyond the regime one can currently test with experiments, GR is generally an ill-behaved theory.
@@
Cosmological observations indicate an expanding Universe whose acceleration can only be explained by introducing ad hoc a cosmological constant, which currently lacks a theoretical explanation and raises several issues (Weinberg, 1989; Peebles & Ratra, 2003). Other observational evidences, such as the rotational curve of galaxies (van Albada et al., 1985; Salucci, 2019) or gravitational lensing effects (Massey et al., 2010) show the presence of a dark massive component of the Universe whose nature is still unknown.
Furthermore, it is well known that GR lacks a quantum description at high-energy scales, and several attempts have been made to create a theory valid at all scales (for a review on the state-of-the-art see, e.g.,  Esposito (2011); Kiefer (2023)).
One way to address these inconsistencies between theory and experiments is to directly modify GR, giving rise to a plethora of possible Extended Theories of Gravity (ETG).
@@
A Yukawa-like interaction emerges quite naturally in the weak field limit of several ETGs, e.g., Scalar-Tensor-Vector theories (Moffat, 2006), massive gravity theories (Visser, 1998; Hinterbichler, 2012), theories in higher dimensions with Kaluza-Klein compactification (Bars & Visser, 1986; Hoyle et al., 2001), massive Brans-Dicke theories (Perivolaropoulos, 2010; Alsing et al., 2012) or $f(R)$ theories (Capozziello et al., 2015). But the so-called fifth-force scenario also appears in some specific models for dark matter (Frieman & Gradwohl, 1991; Gradwohl & Frieman, 1992; Carroll et al., 2009).
@@
Due to the importance that a modification of Newtonian gravity would have, the presence of a Yukawa-like contribution has been repeatedly investigated in the past. The fifth force intensity is well constrained at Solar System scales via the motion of planets (Konopliv et al., 2011; Hees et al., 2014; Bergé, 2017; Will, 2018a; Shankaranarayanan & Johnson, 2022), from the Lunar Laser Ranging experiment (Hofmann & Müller, 2018) and also by making use of the planetary ephemerides (Mariani et al., 2023; Fienga & Minazzoli, 2024). Recent constraints have been obtained from asteroids tracking (Tsai et al., 2023, 2024) and test of the Weak Equivalence Principle (Touboul et al., 2022).
@@
The discovery of orbiting stars around the Galactic Center (GC) (Eckart & Genzel, 1996; Schödel et al., 2002; Ghez et al., 2003; Gillessen et al., 2009b, a; Sabha et al., 2012), all located within one arcsecond distance from the supermassive black hole (SMBH) Sagittarius A (Sgr A), allows one to test GR in a completely different environment from the Solar System.
@@
The importance of looking for a fifth force in the GC lies in the fact that many ETGs that predict a Yukawa-like term also display a screening mechanism that suppresses the fifth force contribution at Solar System scales and prevents its detection - explaining why it would be yet unobserved -, while its effect may be different around SMBHs.
@@
The current constraints on the intensity of a fifth force in the GC come from the analysis of SgrA’s shadow by the Event Horizon Telescope (Vagnozzi et al., 2023), from the measurement of the Schwarzschild precession in S2222 motion (GRAVITY Collaboration, 2020; Jovanović et al., 2023, 2024b, 2024a) and from the analysis of S-stars publicly available (or mock) data (Borka et al., 2013; Capozziello et al., 2014; Borka et al., 2021; Zakharov et al., 2016, 2018; de Martino et al., 2021; Della Monica et al., 2022), also including the presence of a (expected) bulk mass distribution around Sgr A (Jovanović et al., 2021) (for a complete and comprehensive review on astrophysical and theoretical constraints at the GC, see de Laurentis et al. (2023)).
@@
In the context of S-stars, a previous work by Hees et al. (2017) with a full analysis of the S2222 data showed that the intensity of such a contribution cannot exceed $\alpha\sim 0.01$ at scales comparable to S2222-SgrA distance.
@@
In this paper we use the astrometric and spectroscopic measurements of the star S2222 collected at the Very Large Telescope (VLT) by GRAVITY, NACO and SINFONI in order to constrain the intensity of a possible Yukawa correction at the GC. Although we do not expect our results to consistently deviate from the estimates obtained in the aforementioned literature, a complete analysis of S2222 including GRAVITY data, which dominate the $\chi^{2}$ due to their very small uncertainties, is still missing. As will be shown, the precision of the GRAVITY instrument allows us to place a significantly stronger constraint than the previous estimates.
@@ 2 Observations
The set of available data $D$ can be divided as follows:
@@
    • 128 data points collected using both the SHARP camera at the New Technology Telescope between 1992 and 2002 ($\sim$ 10 data points, accuracy $\approx 4\,\rm mas$) and the NACO imager at the VLT between 2002 and 2019 (118 data points, accuracy $\approx 0.5\,\rm mas$);
    • 76 data points collected by GRAVITY at the VLT between 2016 and April 2022 (accuracy $\approx 50\,\rm\mu as$).
@@
    • 102 data points collected by SINFONI at the VLT (100 points) and NIRC2 at Keck (2 points) collected between 2000 and March 2022 (accuracy in good conditions $\approx 10-15\,\rm km/s$).
@@ 3.1 Model
The potential we aim to test has the following form:
@@
where $\alpha$ represents the strength of interaction and $\lambda$ is a scale parameter which depends on the specific theory considered. For example, when new massive fields are included in the theory, $\lambda$ represents the Compton wavelength of the field, which is related to the mass by $m_{\varphi}=h/c\lambda$, where $h$ is the Planck constant.
In comparison to previous work, in this paper the Schwarzschild precession in the S2222 orbit is also included, as it has been detected at $10\sigma$ confidence level by the GRAVITY Collaboration (GRAVITY Collaboration, 2020, 2024).
Although a formal parametrized Post Newtonian (PN) treatment is not possible when a massive field is included in the action (Alsing et al., 2012; Poisson & Will, 2012), one can still derive the equations of motion of a test particle assuming the parametrized PN parameters to be $\gamma=\beta=1$. This latter assumption is valid for ETGs that are indistinguishable from GR at $1$PN order and it is supported by different experimental observations, including at the GC (see, e.g. Will (2018b); Hofmann & Müller (2018); GRAVITY Collaboration (2020); Iorio (2024)). On the other hand, non metric theories of gravity or specific subclasses of f(R) and scalar-tensor theories, with parametrized PN parameters significantly different from unity, are excluded from this analysis (Will, 2014).
The total acceleration felt by the star is:
@@
with $\boldsymbol{r}=r\hat{r}$, $\boldsymbol{v}=\left(\dot{r}\hat{r},r\dot{\theta}\hat{\theta},r\dot{\phi}\sin\theta\hat{\phi}\right)$ and $v=|\boldsymbol{v}|$.
The above expression coincides with the $1$PN acceleration derived in Alves et al. (2024) for the two-body problem in massive Brans-Dicke theory.
Section 3.4 will be devoted to compare Eq. (3) with results developed in the literature when extra massive degrees of freedom are included in the theory.
@@ 3.2 Method
The numerical integration of the equations of motion is performed using a Runge-Kutta $4(5)$ method, further details are reported in Appendix A.
During the fit of the S2222 data the Yukawa length scale $\lambda$ is kept fixed, choosing values between $10^{12}\leq\lambda\leq 10^{15}\,\rm m$, while the intensity $\alpha$ is allowed to vary together with other parameters describing the system.
@@
where $e$ is the eccentricity and $a_{\rm sma}$ the semi major axis of the star S2222, $\Omega_{\rm orb}$, $i_{\rm orb}$ and $\omega_{\rm orb}$ are the three angles used to project the star’s orbital frame into the observer reference frame using the procedure reported in Appendix B.1, $t_{p}$ is the time of pericenter passage, $M$ and $R_{0}$ are the SMBH mass and the GC distance, respectively. The additional parameters $\{x_{0},y_{0},v_{x_{0}},v_{y_{0}},v_{z_{0}}\}$ characterize the NACO/SINFONI data reference frame with respect to Sgr A (Plewa et al., 2015).
To fit S2222 data, we perform a Markov Chain Monte Carlo (MCMC) analysis using the Python package emcee (Foreman-Mackey et al., 2013). The log-likelihood is given by
@@
The priors are listed in Tables  1 and 2. Uniform priors are used for the physical parameters, that is, we only imposed physically motivated bounds, while Gaussian priors are implemented for the offset parameters, since the latter have been well constrained by an independent previous work and are not expected to change (Plewa et al., 2015). The initial points $\Theta_{i}^{0}$ in the MCMC are chosen to be the most recent best fit parameters of S2222 orbit reported in the literature (GRAVITY Collaboration, 2024).
@@
Table 2: Gaussian priors used in the MCMC analysis. $\xi$ and $\sigma$ represent the mean and the standard deviation of the Gaussian distributions, respectively, and they come from Plewa et al. (2015).
@@
$v_{z_{0}}\,\rm[km/s]$2.455 0 5
In the sampling phase of the MCMC implementation, we used 64 walkers and $10^{5}$ iterations. The burning-in phase is skipped and the last $80\%$ of the chains is used to compute the mean and standard deviation of the posterior distributions of the parameters. The convergence of the MCMC analysis is ensured by means of the autocorrelation time $\tau_{c}$, that is, we ran $N$ iterations such that $N\gg 50\,\tau_{c}$.
@@ 3.3 Results
One can classify three different regimes in the posterior distributions $P(|\alpha||D)$, according to the value of $\lambda$ with respect to the orbital range of S2222, which is $1.7\cdot 10^{13}\,\rm m\lesssim r_{s2}\lesssim 1.5\cdot 10^{14}\,\rm m$.
@@
When $\lambda\ll r_{s2}$, the acceleration is no longer dependent on the parameter $\alpha$ and no meaningful constraints can be obtained in this regime. The small difference in the $95\%$ upper limit on $|\alpha|$ with the UCLA group resides in the different model implemented to fit the data, i.e., Hees et al. (2017) considered no Schwarzschild precession but rather an extended mass with power law distribution, which is absent in our work. The reason behind this choice comes from the fact that the presence of a spherically symmetric mass distribution around SgrA has been extensively tested by the GRAVITY Collaboration (GRAVITY Collaboration, 2018, 2022, 2024), finding a stringent upper limit of $M_{\rm ext}\lesssim 1200\,M_{\odot}$. Taking into account this current upper limit, we decided to neglect the presence of an extended mass, as it would not alter our results.
@@
When $\lambda\sim r_{s2}$, the best constraints on $|\alpha|$ are obtained, finding the most stringent upper limit $|\alpha|<0.003$ for $\lambda=3\cdot 10^{13}\,\rm m\sim 200\,\rm AU$. This limit improves the previous estimate of Hees et al. (2017), which reported $|\alpha|<0.016$ for $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,\rm m$, stressing the importance of the precision of the GRAVITY instrument.
@@
Finally, when $\lambda\gg r_{s2}$, the only component left in the equations of motion is the monopolar term $M(1+\alpha)/r$. This corresponds to a simple rescaling of the mass term, and hence in this regime $M$ and $\alpha$ are completely degenerate and they can not be constrained separately. To obtain the upper limit on $\alpha$ in this regime, the bounds on $M$ in Table 1 have been extended to $M\in(10^{-4},10^{4})\cdot 10^{6}\,M_{\odot}$, and the same bounds have been used for $\alpha$.
@@
A summary of the above results is reported in Figure 2, where the $95\%$ confidence interval on $|\alpha|$ as function of $\lambda$ is shown. Those confidence intervals are estimated as $3$ times the standard deviation when the posterior distributions are Normal or by computing the upper limit that corresponds to $95\%$ of the area below the curve when the distributions have different shapes.
We note that GRAVITY data produce an overall improvement of roughly one order of magnitude over the entire parameter space tested.
@@
If one assumes that the gravitational interaction is mediated by a massive boson as in massive gravity theories (where $\alpha=1$), the length scale $\lambda$ corresponds to the Compton wavelength of the particle and hence an upper limit on the graviton’s mass can be derived. Since $\alpha=1$ is excluded at $95\%$ confidence level for $\lambda\lesssim 8\cdot 10^{14}\,\rm m$, this lower bound on the wavelength $\lambda$ can be translated into an upper limit on the graviton mass, corresponding to $m_{g}\lesssim 2.5\cdot 10^{-22}\,\rm eV$.
@@
In the regime $\lambda\sim r_{s2}$, when no $1$PN acceleration is included, the presence of the Yukawa term in the equation of motion induces a prograde precession comparable to the Schwarzschild one. This is shown in Figure 1 for $\lambda=10^{13}\,\rm m$, where one can see that the posterior distribution $P(\alpha|D)$ is a Gaussian with mean around $\alpha\sim 0.0026$, as opposed to the $1$PN posterior.
@@
Following Adkins & McDonnell (2007) one can compute the precession angle induced by a potential in a full orbit as
@@
where $U(z)$ is the perturbing potential evaluated at radius $r=L/(1+ez)$ with $L=a_{\rm sma}(1-e^{2})$. For $\alpha=0.0026$, this corresponds to $\Delta\phi_{p}\sim 0.13^{\circ}$. Taking the most up-to-date value of the Schwarzschild precession reported in GRAVITY Collaboration (2024), $\Delta\phi_{\rm Sch}=12.1^{\prime}\times(0.911\pm 0.13)=(0.18\pm 0.03)^{\circ}$, one can see that the precession angle induced by the Yukawa potential is compatible, within the $2\sigma$ uncertainties, with $\Delta\phi_{\rm Sch}$.
Figure 1: Comparison of the posterior distributions $P(|\alpha||D)$ between the Keplerian model (black curve) and the $1$PN model (red curve).
Figure 2: $95\%$ confidence level on $|\alpha|$ obtained in this work (red dots) compared with previous estimates by Hees et al. (2017) (blue dots). The dotted vertical line represents the minimum of the curve, in correspondence of $\lambda\sim 3\cdot 10^{13}\,\text{m}\sim 200\,\text{AU}$

### v4

# Exploring the presence of a fifth force at the Galactic Center

@@ Abstract
Aims. We investigate the presence of a Yukawa-like correction to Newtonian gravity at the Galactic Center, leading to a new upper limit on the intensity of such a correction.
Methods. We performed a Markov chain Monte Carlo (MCMC) analysis using the astrometric and spectroscopic data of star S22 collected at the Very Large Telescope by GRAVITY, NACO, and SINFONI instruments, covering the period from $1992$ to $2022$.
Results. The precision of the GRAVITY instrument allows us to derive the most stringent upper limit at the Galactic Center for the intensity of the Yukawa contribution ($\propto\,\alpha e^{-\lambda r}$) of $|\alpha|<0.003$ for a scale length of $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,\rm AU)$. This is an improvement on all estimates obtained in previous works by roughly one order of magnitude.
@@ 1 Introduction
General relativity (GR) is the most widely recognized theory of gravity today. Its predictions have been extensively tested on Solar System scales and using gravitational waves emission by black holes (BHs) and binary pulsars (Will, 2014, 2018b; Nitz et al., 2021). Until now, no significant deviation from GR has been detected in any of these observations. However, it is also known that beyond the regime we are currently able to test experimentally, GR is generally considered insufficient.
@@
Cosmological observations indicate an expanding Universe whose acceleration can only be explained by introducing a cosmological constant ad hoc, which lacks a solid theoretical explanation and raises several issues (Weinberg, 1989; Peebles and Ratra, 2003). Other observational evidence, such as the rotational curve of galaxies (van Albada et al., 1985; Salucci, 2019) or gravitational lensing effects (Massey et al., 2010) indicate the presence of a dark massive component of the Universe, whose nature is still unknown. Furthermore, it is well known that GR lacks a quantum description at high-energy scales and several attempts have been made to create a theory valid at all scales (for a review on the state-of-the-art approaches, see, e.g., Esposito (2011); Kiefer (2023)).
@@
One way to address these inconsistencies between theory and experiments is to directly modify GR, giving rise to a plethora of possible extended theories of gravity (ETG). In particular, a Yukawa-like interaction emerges quite naturally in the weak field limit of several ETGs; for instance, scalar-tensor-vector theories (Moffat, 2006), massive gravity theories (Visser, 1998; Hinterbichler, 2012), theories in higher dimensions with Kaluza-Klein compactification (Bars and Visser, 1986; Hoyle et al., 2001), massive Brans-Dicke theories (Perivolaropoulos, 2010; Alsing et al., 2012), or $f(R)$ theories (Capozziello et al., 2015). However, the so-called ”fifth-force” scenario has also made an appearance in a number of specific dark matter models (Frieman and Gradwohl, 1991; Gradwohl and Frieman, 1992; Carroll et al., 2009).
@@
Due to the importance that a modification of Newtonian gravity would have on our understanding of the Universe, the presence of a Yukawa-like contribution has been repeatedly investigated in the past. The fifth-force intensity is well constrained at Solar System scales via the motion of planets (Konopliv et al., 2011; Hees et al., 2014; Bergé, 2017; Will, 2018a; Shankaranarayanan and Johnson, 2022), from the Lunar Laser Ranging experiment (Hofmann and Müller, 2018) and also by making use of the planetary ephemerides (Mariani et al., 2023; Fienga and Minazzoli, 2024). Recent constraints have been obtained from asteroids tracking (Tsai et al., 2023, 2024) and test of the weak equivalence principle (Touboul and others, 2022).
@@
The discovery of orbiting stars around the Galactic Center (GC) (Eckart and Genzel, 1996; Schödel et al., 2002; Ghez et al., 2003; Gillessen et al., 2009b, a; Sabha and others, 2012), all located within one arcsecond distance from the supermassive black hole (SMBH) Sagittarius A (Sgr A), allows us to test GR in a completely different environment from the Solar System.
@@
The importance of looking for a fifth force in the GC lies in the fact that many ETGs that predict a Yukawa-like term also display a screening mechanism that suppresses the fifth force contribution at Solar System scales and prevents its detection. This would explain why it would be yet unobserved, while its effect may be different around SMBHs.
@@
The current constraints on the intensity of a fifth force in the GC come from the analysis of SgrA’s shadow by the Event Horizon Telescope (Vagnozzi and others, 2023), from the measurement of the Schwarzschild precession in S22 motion (GRAVITY Collaboration, 2020; Jovanović et al., 2023, 2024b, 2024a) and from the analysis of S-stars publicly available (or mock) data (Borka et al., 2013; Capozziello et al., 2014; Borka et al., 2021; Zakharov et al., 2016, 2018; de Martino et al., 2021; Della Monica et al., 2022), while also including the presence of a (expected) bulk mass distribution around Sgr A (Jovanović et al., 2021). For a complete and comprehensive review on astrophysical and theoretical constraints at the GC, we refer to de Laurentis et al. (2023).
@@
In the context of S-stars, a previous work by Hees and others (2017), with a full analysis of the S22 data, showed that the intensity of such a contribution cannot exceed $\alpha\sim 0.01$ at scales comparable to the S22-SgrA distance.
@@
In this Letter, we use the astrometric and spectroscopic measurements of the star S22 collected at the Very Large Telescope (VLT) by GRAVITY, NACO, and SINFONI to constrain the intensity of a possible Yukawa correction at the GC. Although we do not expect our results to consistently deviate from the estimates obtained in the aforementioned literature, a complete analysis of S22 including GRAVITY data, which dominate the $\chi^{2}$ due to their very small uncertainties, is still lacking. As we show here, the precision of the GRAVITY instrument allows us to place a significantly stronger constraint than the previous estimates.
@@ 2 Observations
The set of available data, $D,$ can be organized according to the following criteria:
@@
    • 128 data points collected using both the SHARP camera at the New Technology Telescope between 1992 and 2002 ($\sim$ 10 data points, accuracy of $\approx 4\,\rm mas$) and the NACO imager at the VLT between 2002 and 2019 (118 data points, accuracy of $\approx 0.5\,\rm mas$);
    • 76 data points collected by GRAVITY at the VLT between 2016 and April 2022 (accuracy of $\approx 50\,\rm\mu as$).
@@
    • 102 data points collected by SINFONI at the VLT (100 points) and NIRC2 at Keck (2 points) collected between 2000 and March 2022 (accuracy in good conditions of $\approx 10-15\,\rm km/s$).
@@ 3.1 Model
The potential we aim to test takes the following form:
@@
where $\alpha$ represents the strength of interaction and $\lambda$ is a scale parameter, which depends on the specific theory considered. For example, when new massive fields are included in the theory, $\lambda$ represents the Compton wavelength of the field, which is related to the mass by $m_{\varphi}=h/c\lambda$, where $h$ is the Planck constant.
In comparison to previous works, here the Schwarzschild precession in the S22 orbit is also included, since it has been formally detected at a $10\sigma$ confidence level by the GRAVITY Collaboration (GRAVITY Collaboration, 2020, 2024).
Although a formal parametrized post-Newtonian (PN) treatment is not possible when a massive field is included in the action (Alsing et al., 2012; Poisson and Will, 2012), we can still derive the equations of motion of a test particle, assuming the parametrized PN parameters to be $\gamma=\beta=1$. This latter assumption is valid for ETGs that are indistinguishable from GR at 11PN order and it is supported by different experimental observations, including those carried out for the GC (see, e.g., Will (2018b); Hofmann and Müller (2018); GRAVITY Collaboration (2020); Iorio (2024)). On the other hand, non-metric theories of gravity or specific subclasses of f(R) and scalar-tensor theories, with parametrized PN parameters that significantly stray from unity, are excluded from this analysis (Will, 2014).
The total acceleration experienced by the star is
@@
[…]
The above expression coincides with the 11PN acceleration derived in Alves et al. (2024) for the two-body problem in massive Brans-Dicke theory. Section 3.4 will be devoted to a comparison of Eq. (3) with results developed in the literature when extra massive degrees of freedom are included in the theory.
@@ 3.2 Method
The numerical integration of the equations of motion was performed using a Runge-Kutta $4(5)$ method, with further details reported in Appendix A. During the fit of the S22 data, the Yukawa length scale, $\lambda$, was kept fixed, with values set between $10^{12}\leq\lambda\leq 10^{15}\,\rm m$, while the intensity, $\alpha$, was allowed to vary together with other parameters describing the system.
@@
where $e$ is the eccentricity and $a_{\rm sma}$ the semi major axis of the star S22, while $\Omega_{\rm orb}$, $i_{\rm orb}$, and $\omega_{\rm orb}$ are the three angles used to project the star’s orbital frame into the observer reference frame using the procedure reported in Appendix B.1. Then, $t_{p}$ is the time of pericenter passage, while $M$ and $R_{0}$ are the SMBH mass and the GC distance, respectively. The additional parameters $\{x_{0},y_{0},v_{x_{0}},v_{y_{0}},v_{z_{0}}\}$ characterize the NACO/SINFONI data reference frame with respect to Sgr A (Plewa et al., 2015).
To fit the S22 data, we perform a Markov chain Monte Carlo (MCMC) analysis using the Python package emcee (Foreman-Mackey et al., 2013). The log-likelihood is given by
@@
The priors are listed in Tables 1 and 1. Uniform priors were used for the physical parameters, that is, we only imposed physically motivated bounds, while Gaussian priors were implemented for the offset parameters, since the latter have been well constrained by an independent previous work and were not expected to change (Plewa et al., 2015). The initial points $\Theta_{i}^{0}$ in the MCMC are chosen to be the most recent best fit parameters of S22 orbit reported in the literature (GRAVITY Collaboration, 2024).
@@
[…]
@@
$v_{z_{0}}\,\rm[km/s]$ -2.455 0 5
In the sampling phase of the MCMC implementation, we used 64 walkers and $10^{5}$ iterations. The burning-in phase was skipped and the last $80\%$ of the chains was used to compute the mean and standard deviation of the posterior distributions of the parameters. The convergence of the MCMC analysis was ensured by means of the autocorrelation time $\tau_{c}$, that is, we ran $N$ iterations such that $N\gg 50\,\tau_{c}$.
@@ 3.3 Results
We can classify three different regimes in the posterior distributions $P(|\alpha||D)$, according to the value of $\lambda$ with respect to the orbital range of S22, which is $1.7\cdot 10^{13}\,\rm m\lesssim r_{s2}\lesssim 1.5\cdot 10^{14}\,\rm m$. When $\lambda\ll r_{s2}$, the acceleration is no longer dependent on the parameter $\alpha$ and no meaningful constraints can be obtained in this regime. The small difference in the $95\%$ upper limit on $|\alpha|$ with the UCLA group resides in the different model implemented to fit the data.
@@
In particular, Hees and others (2017) considered no Schwarzschild precession but an extended mass with power law distribution instead, which is absent from our work. The reason behind this choice comes from the fact that the presence of a spherically symmetric mass distribution around SgrA has been extensively tested by the GRAVITY Collaboration (GRAVITY Collaboration, 2018, 2022, 2024), finding a stringent upper limit of $M_{\rm ext}\lesssim 1200\,M_{\odot}$. Taking into account this current upper limit, we decided to neglect the presence of an extended mass, as it would not alter our results.
@@
When $\lambda\sim r_{s2}$, the best constraints on $|\alpha|$ can be obtained, determining the most stringent upper limit of $|\alpha|<0.003$ for $\lambda=3\cdot 10^{13}\,\rm m\sim 200\,\rm AU$. This limit improves the previous estimate of Hees and others (2017), which reported $|\alpha|<0.016$ for $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,\rm m$, stressing the importance of the precision of the GRAVITY instrument.
@@
Finally, when $\lambda\gg r_{s2}$, the only component left in the equations of motion is the monopolar term $M(1+\alpha)/r$. This corresponds to a simple rescaling of the mass term and, hence, in this regime $M$ and $\alpha$ are completely degenerate and they cannot be constrained separately. To obtain the upper limit on $\alpha$ in this regime, the bounds on $M$ in Table 1 have been extended to $M\in(10^{-4},10^{4})\cdot 10^{6}\,M_{\odot}$, and the same bounds have been used for $\alpha$.
@@
A summary of the above results is reported in Figure 2, where the $95\%$ confidence interval on $|\alpha|$ as function of $\lambda$ is shown. Those confidence intervals are estimated as three times the standard deviation when the posterior distributions are normal or by computing the upper limit that corresponds to $95\%$ of the area below the curve when the distributions have different shapes. We note that GRAVITY data produce an overall improvement of roughly one order of magnitude over the entire parameter space tested.
@@
If we assume that the gravitational interaction is mediated by a massive boson as in massive gravity theories (where $\alpha=1$), the length scale $\lambda$ corresponds to the Compton wavelength of the particle; hence, an upper limit on the graviton’s mass can be derived. Since $\alpha=1$ is excluded at $95\%$ confidence level for $\lambda\lesssim 8\cdot 10^{14}\,\rm m$, this lower bound on the wavelength, $\lambda$, can be translated into an upper limit on the graviton mass, corresponding to $m_{g}\lesssim 2.5\cdot 10^{-22}\,\rm eV$.
@@
In the regime where $\lambda\sim r_{s2}$, when no 11PN acceleration is included, the presence of the Yukawa term in the equation of motion induces a prograde precession comparable to the Schwarzschild one. This is shown in Figure 1 for $\lambda=10^{13}\,\rm m$, where we can see that the posterior distribution $P(\alpha|D)$ is a Gaussian with mean around $\alpha\sim 0.0026$, as opposed to the 11PN posterior.
@@
Following Adkins and McDonnell (2007), we can compute the precession angle induced by a potential in a full orbit as
@@
where $U(z)$ is the perturbing potential evaluated at radius $r=L/(1+ez)$ with $L=a_{\rm sma}(1-e^{2})$. For $\alpha=0.0026$, this corresponds to $\Delta\phi_{p}\sim 0.13^{\circ}$. Taking the most up-to-date value of the Schwarzschild precession reported in GRAVITY Collaboration (2024), $\Delta\phi_{\rm Sch}=12.1^{\prime}\times(0.911\pm 0.13)=(0.18\pm 0.03)^{\circ}$, we can see that the precession angle induced by the Yukawa potential is compatible, within the $2\sigma$ uncertainties, with $\Delta\phi_{\rm Sch}$.
Figure 1: Comparison of the posterior distributions $P(|\alpha||D)$ between the Keplerian model (black curve) and the 11PN model (red curve).
Figure 2: $95\%$ confidence level on $|\alpha|$ obtained in this work (red dots), compared with previous estimates by Hees and others (2017) (blue dots). The dotted vertical line represents the minimum of the curve, in correspondence to $\lambda\sim 3\cdot 10^{13}\,\text{m}\sim 200\,\text{AU}$